OpenStudy (anonymous):

Am I on the right track? 2x - y = 8, y = 8 -2x

OpenStudy (anonymous):

yess

OpenStudy (phi):

are you trying to solve for y? start with 2x-y= 8 add y to both sides then add -8 to both sides

OpenStudy (anonymous):

what is the next step?

OpenStudy (phi):

2x-y=8 add y to both sides 2x-y+y= 8+y can you simplify the left hand side?

OpenStudy (anonymous):

yeah 2x-2y=8+y

OpenStudy (phi):

? -y+y = -2y????

OpenStudy (anonymous):

oh oops hahaha I thought there were 2 positive y's sorry. No they just cancel out right?

OpenStudy (phi):

yes

OpenStudy (phi):

now add -8 to both sides

OpenStudy (anonymous):

ok 2x=8+y then it would become 2x=y

OpenStudy (phi):

now add -8 to BOTH sides

OpenStudy (phi):

we have to keep things in balance.

OpenStudy (anonymous):

sorry sorry. 2x=8+y to 2x-8=y

OpenStudy (phi):

yes, and you can swap the order 2x-8=y is the same as y= 2x-8 (in case have get a multiple choice)

OpenStudy (anonymous):

ok :D thank you!!!! You are a very patient teacher.